Proof (a+b) 2 in full detail. Please answer it as soon as possible. Thankyou.

Proof (a+b)2 in full detail.
 
 
Please answer it as soon as possible. 
 
Thankyou.

Grade:11

2 Answers

Arun
5 years ago

.Just do it right away. 

(a+b)^2 = (a+b) * ( a+b) 
(a+b) * (a+b) multply each element indevidually 

= aa+ab+ba+bb = a2+2ab+b2

Ahmed
5 years ago
(a+b)2=(a+b)(a+b). 
            =ab+ab+a2+b2.
 (a+b)2 = a2 + b +2ab
 
 
Hence proved.
 
 
 
 
Thankyou..................................